OpenStudy (anonymous):

Find a polynomial fx of degree 4 that has the following zeros. −5 , 0 , 3 , −2

OpenStudy (jdoe0001):

\( -5 , 0 , 3 , -2\quad \begin{cases} x=-5\to x+5=0\to &(x+5)=0\\ x=0\to x-0=0\to &(x)=0\\ x=3\to x-3=0\to &(x-3)=0\\ x=-2\to x+2=0\to &(x+2)=0 \end{cases} \\ \quad \\ \bf (x+5)(x)(x-3)(x+2)=\textit{original polynomial} \)